# INSIGHT: Universal Neural Simulator for Analog Circuits Harnessing Autoregressive Transformers

### v1

@@ 3.2.1   INSIGHT Pre-Training
As depicted in the left part of Figure 2, the pre-training phase involves a one-time effort to generate a simulation-based performance dataset, followed by a one-time training of our proposed neural simulator, INSIGHT. The data collection involves gathering a random dataset of Gaussian-sampled design parameters and their corresponding performance metrics derived from SPICE simulations. To address computational demands, the initial dataset size is defined as a hyperparameter, allowing for adjustments tailored to specific use cases.
This flexibility helps balance the resource usage during the initial data collection with the subsequent sample efficiency following a user query. This process is followed by the one-time pre-training of our proposed neural simulator, INSIGHT. By employing transformer technology, INSIGHT effectively utilizes transfer learning, thereby reducing dataset requirements across different technologies.
3.3. INSIGHT-M
The pre-trained INSIGHT is subsequently integrated into a Model-Based Batch RL framework for circuit sizing, referred to as INSIGHT-M. As shown in the right part of Figure 2, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance target specifications, including previously unseen ones. In our experiments, we utilized a Proximal Policy Optimization (PPO)-based RL agent 44, a widely used approach in Reinforcement Learning. PPO has been extensively utilized in analog design optimization, as in  45, 24. PPO operates by executing multiple trajectory rollouts in parallel, followed by their subsequent aggregation to update the policy under certain constraints. Although PPO’s use of multiple trajectories under safe constraints enhances the robustness and safety of the learning process, it tends to compromise sample efficiency. The INSIGHT-M framework addresses this drawback by enabling the RL agents to be pre-trained within the close-to-real synthetic neural simulator environment, INSIGHT, until they achieve satisfactory performance. Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy. This process involves periodic performance evaluations in the real environment. If the agents perform adequately, their deployment persists. If not, new real-world data is acquired to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated. This iterative refinement process optimizes the balance between safety, robustness, and efficiency.
Central to the efficiency of INSIGHT-M is the training of a robust RL agent within a synthetic environment that closely mimics real-world conditions. The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency. Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, enhances the agent’s adaptability to real-world environments. Furthermore, the framework’s capability for iterative refinement effectively balances exploration and exploitation, achieving high sample efficiency.
4. Experiments
@@
Figure 3. Benchmark Circuits: (a) Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, (b) Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, (c) Two-Stage Transimpedance Amplifier, (d) Three-Stage Transimpedance Amplifier, (e) Comparator, (f) Level Shifter.
@@
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¿0.99 0.65 ¿0.99 0.62
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¿0.99 0.73 ¿0.99 0.77
Two-Stage Transimpedance Amplifier ¿0.99 0.81 ¿0.99 0.78
Three-Stage Transimpedance Amplifier ¿0.99 0.91 ¿0.99 0.88
Comparator 0.70 0.60 0.97 0.61
Level Shifter ¿0.99 ¡0.00 0.91 0.07
@@
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair ¡0.001 0.01 ¡0.001 0.01
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair ¡0.001 0.05 ¡0.001 0.04
Two-Stage Transimpedance Amplifier ¡0.001 0.05 ¡0.001 0.05
Three-Stage Transimpedance Amplifier ¡0.001 0.01 ¡0.001 0.01
Comparator 0.010 0.01 ¡0.001 0.01
Level Shifter ¡0.001 0.03 ¡0.001 0.05
@@
Circuit Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Two-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
Three-Stage Transimpedance Amplifier 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001 1500-500 ¿0.99 ¡0.001
3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001 3000-1000 ¿0.99 ¡0.001
@@
Algorithm INSIGHT-M (with PPO) AutoCkt
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1,500 steps 14,100 steps
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1,500 steps 9,300 steps
Two-Stage Transimpedance Amplifier 300 steps 11,100 steps
Three-Stage Transimpedance Amplifier 600 steps 8,700 steps
Comparator 1,200 steps 8,400 steps
4.1. Benchmark Tasks and Algorithms
In this paper, we utilized a GPU machine with x8 NVIDIA P100 GPUs under Intel(R) Xeon(R) CPU E5-2698 V4 2.220GHz for training and inference. We used the following benchmark circuits to evaluate INSIGHT and to validate our approach: Two-Stage Miller-Compensated OPAMP with NMOS Input Pair, Two-Stage Miller-Compensated OPAMP with PMOS Input Pair, Two-Stage Transimpedance Amplifier, Three-Stage Transimpedance Amplifier, Comparator, and Level Shifter. Figure 3 illustrates the detailed topology of the benchmark circuits. Table 6 presents the performance specifications for each circuit we focussed on for this paper. We have validated our concept using BSIM 45nm, Skywater 130nm, BSIM 180nm, and FreePDK 45nm technology nodes.
@@ 4.2.1   Prediction Accuracy
Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  27, which serves as the baseline for our comparison with INSIGHT. Table 3 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on the BSIM 45nm technology node. As demonstrated in the table, our analysis for the comparator and level shifter cases involved a combination of DC and simulation-costly transient specifications. Specifically for the level shifter, leveraging its autoregressive capabilities, INSIGHT enhances the predictability of simulation-intensive transient specifications such as Avg. Delay and Delay Balance by efficiently utilizing economical metric information like DC Power and Ratio. The table further details INSIGHT’s performance across various training and testing data sizes, indicating robust performance for dataset sizes close to 1500-1800. Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times. In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases. Table 4 validates INSIGHT’s effectiveness across 45nm, 130nm, and 180nm technology nodes, reinforcing its universal applicability. Our experiments showed that simple transfer learning or knowledge-based domain adaptation techniques can reduce initial dataset requirements.
@@ 4.3. INSIGHT-M Performance
We evaluate the performance of INSIGHT-M against AutoCkt 24, an established online learning-based sizing algorithm based on Proximal Policy Optimization (PPO), using default hyperparameter settings. Table 5 highlights the performance comparison. Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance. For these results, the INSIGHT model used within INSIGHT-M was pre-trained on a minimal initial dataset of 300 samples, showcasing the efficacy of our neural simulator INSIGHT in optimization frameworks.
Table 6. Performance Specifications for each circuit
Circuit Performance Specifications Width (Capacitor) Range ($\mu$m ($pF$))
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair $I_{Q}$ (mA) DC Gain (dB) UGBW (MHz) Phase Margin (degrees) [1.0e-1, 3.0e1]
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair [1.0e-1, 3.0e1]
Two-Stage Transimpedance Amplifier [1.0e-1, 1.0e2]
Three-Stage Transimpedance Amplifier [1.0e-2, 3.0e1]
Comparator DC Power (W) Avg. Delay (sec) [1.0e-1, 3.0e1]
Level Shifter DC Power (W) Ratio Avg. Delay (sec) Delay Balance (sec) [4.6e-1, 1.0e2]
@@ 5. Conclusion
This paper proposes INSIGHT, a novel, GPU-powered, technology-independent, effective universal neural simulator in the analog front-end automation loop. Leveraging self-attention to understand underlying inherent parameter-performance interdependencies, INSIGHT accurately predicts performance across analog circuits and their corresponding metrics over various technology nodes, achieving inference times within microseconds. Notably, its autoregressive capabilities enable accurate prediction of simulation-costly critical transient specifications leveraging less expensive performance metric information. Due to its low cost and high fidelity, INSIGHT serves as an effective substitute for standard simulators in analog front-end optimization frameworks, facilitating effective and efficient design space exploration and improved sample efficiency. Our Model-Based Batch RL framework, INSIGHT-M, which utilizes INSIGHT for analog sizing, achieves at least 50×50$\times$50 × improvement in sample efficiency across circuits. We rigorously validated our approach on diverse circuit types across technology nodes. To our knowledge, this marks the first application of autoregressive transformers in analog front-end design automation. We plan to extend this technology to address more complex analog design challenges, including mismatch analysis and design migration. We aim to enhance our adaptation techniques to mitigate synthetic to real-world adaptation errors, thereby strengthening the robustness and reliability of our models under dynamic conditions. This innovative application of autoregressive transformers opens new avenues for leveraging advanced NLP techniques, including grammatical models and language-based inversion techniques in circuit modeling and optimization, inspiring further innovation in the field.

### v2

@@ 3.2.1   INSIGHT Pre-Training
As depicted in the left part of Figure 2, the pre-training phase involves a one-time effort to generate a simulation-based performance dataset, followed by a one-time training of our proposed neural simulator, INSIGHT. The data collection involves gathering a random dataset of Gaussian-sampled design parameters and their corresponding performance metrics derived from SPICE simulations. To address computational demands, the initial dataset size is defined as a hyperparameter, allowing for adjustments tailored to specific use cases. This flexibility helps balance the resource usage during the initial data collection with the subsequent sample efficiency following a user query. Additionally, INSIGHT can leverage transfer learning to reduce dataset requirements across technologies while maintaining high performance.
@@
Figure 3. Benchmark Circuits: (a) 2-Stage Miller-Compensated OPAMP with NMOS Input Pair, (b) 2-Stage Miller-Compensated OPAMP with PMOS Input Pair, (c) 2-Stage TIA, (d) 3-Stage TIA, (e) Comparator, (f) Level Shifter.
Table 3. Performance Metrics for each circuit
Circuits Performance Metrics
2-Stage Miller-Compensated OPAMP with NMOS Input Pair $I_{Q}$ (mA) DC Gain (dB) UGBW (MHz) Phase Margin (${}^{\circ})$
2-Stage Miller-Compensated OPAMP with PMOS Input Pair
2-Stage TIA
3-Stage TIA
Comparator DC Power (W) Avg. Delay (sec)
Level Shifter DC Power (W) Ratio Avg. Delay (sec) Delay Balance (sec)
3.2.2   INSIGHT for Optimization
The pre-trained INSIGHT is subsequently integrated into a Model-based Batch RL framework for circuit sizing, referred to as INSIGHT-M. As shown in the right part of Figure 2, upon query, INSIGHT-M is deployed to optimize device sizing to meet various user-specified performance targets. In our experiments, we use PPO-based RL agents  35 with default hyperparameters, a method extensively used in EDA because of its robustness and effectiveness  36, 17. PPO enhances learning robustness and safety by executing multiple parallel trajectory rollouts and aggregating them to update policies under safe constraints but at the expense of poor sample efficiency. The INSIGHT-M framework addresses this drawback by training robust RL agents within INSIGHT’s synthetic environment, which closely mimics real-world conditions. Once trained, these agents are deployed in the real-world simulator environment using a batch RL strategy. This process involves periodic performance evaluations in the real environment. If the agents perform adequately, their deployment persists. If not, newly acquired real-world data combined with efficiently sampled previously collected data is utilized to fine-tune INSIGHT, and the cycle of evaluation and adjustment is repeated. This iterative refinement process optimizes the balance between safety, robustness, and efficiency. The high fidelity of INSIGHT ensures that the agent’s learning and subsequent behaviors are closely aligned with real-world design tasks, thereby minimizing trial and error and improving sample efficiency. Incorporating robust adaptation techniques, such as uncertainty-driven Upper Confidence Bound (UCB) based exploration, further enhances the agent’s adaptability to real-world environments. Please note that other RL algorithms can be employed with INSIGHT.
4. Experiments
@@
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 0.9901 0.6578 0.9960 0.6240
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 0.9937 0.7270 0.9973 0.7872
Two-Stage Transimpedance Amplifier 0.9982 0.8020 0.9933 0.7784
Three-Stage Transimpedance Amplifier 0.9997 0.9290 0.9996 0.8894
Comparator 0.7231 0.5537 0.998 0.6535
Level Shifter 0.9905 -0.1554 0.9265 0.1252
@@
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 0.00053 0.02054 0.00054 0.01924
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 0.00023 0.04267 0.00037 0.04267
Two-Stage Transimpedance Amplifier 0.00040 0.04335 0.00024 0.04475
Three-Stage Transimpedance Amplifier 0.00012 0.01688 0.00018 0.01727
Comparator 0.00666 0.01856 0.00160 0.01856
Level Shifter 0.00160 0.04436 0.00047 0.05950
@@
Circuits Train Test Size R2 MSE Train Test Size R2 MSE Train Test Size R2 MSE
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1500-500 0.9901 0.00053 1500-500 0.9988 0.00027 1500-500 0.9961 0.00046
3000-1000 0.9960 0.00054 3000-1000 0.9990 0.00082 3000-1000 0.9995 0.00094
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1500-500 0.9937 0.00023 1500-500 0.9982 0.00041 1500-500 0.9971 0.00011
3000-1000 0.9973 0.00037 3000-1000 0.9951 0.00010 3000-1000 0.9961 0.00019
Two-Stage Transimpedance Amplifier 1500-500 0.9982 0.00040 1500-500 0.9975 0.00065 1500-500 0.9901 0.00050
3000-1000 0.9933 0.00024 3000-1000 0.9991 0.00023 3000-1000 0.9980 0.00047
Three-Stage Transimpedance Amplifier 1500-500 0.9997 0.00012 1500-500 0.9956 0.00085 1500-500 0.9969 0.00011
3000-1000 0.9996 0.00018 3000-1000 0.9977 0.00011 3000-1000 0.9955 0.00019
@@
Algorithm INSIGHT-M (with PPO) AutoCkt (PPO)
Circuits Pre-training Data Real-time Simulations Real-time Simulations
Two-Stage Miller-Compensated OPAMP with NMOS Input Pair 1,500 $2\pm 0.18$ steps 14,100 steps
Two-Stage Miller-Compensated OPAMP with PMOS Input Pair 1,500 $2\pm 0.18$ steps 9,300 steps
Two-Stage Transimpedance Amplifier 300 $15\pm 5$ steps 11,100 steps
Three-Stage Transimpedance Amplifier 600 $15\pm 5$ steps 8,700 steps
Comparator 1,200 $15\pm 5$ steps 8,400 steps
4.1. Benchmark Circuits and Technologies
We trained and evaluated INSIGHT on a machine with eight NVIDIA P100 GPUs and an Intel Xeon E5-2698 V4 CPU, using the following benchmark circuits to validate our approach: 2-Stage Miller-Compensated OPAMP with NMOS Input Pair, 2-Stage Miller-Compensated OPAMP with PMOS Input Pair, 2-Stage TIA (Transimpedance Amplifier), 3-Stage TIA, Comparator, and Level Shifter. Figure 3 illustrates the detailed topology of the benchmark circuits. Table 3 lists the performance metrics for each circuit we focussed on in this paper. We have validated our concept using BSIM 45nm, Skywater 130nm, BSIM 180nm, and FreePDK 45nm technologies.
@@ 4.2.1   Prediction Accuracy
Table 2 provides the architectural details for the FC Ensemble-based neural simulator used in CRONuS  25, which serves as the baseline for our comparison with INSIGHT. We also validate INSIGHT across various types of performance metrics. As illustrated in Table 3, different combinations of economic metrics and simulation-intensive transient specifications have been considered for the comparator and level shifter circuits. Table 4 demonstrates that INSIGHT consistently achieves high R2 scores and low MSE losses across all benchmark circuits on 45nm technologies. The table further details INSIGHT’s performance across various training dataset sizes, indicating robust performance for dataset sizes $\sim$ 1500-1800. Given that simulation data collection is a one-time effort, this investment can prove cost-effective if a topology is to be optimized multiple times. In our optimization framework, we define the initial dataset size as a hyperparameter to efficiently balance computational resource utilization with subsequent sample efficiency, allowing for adjustments tailored to specific use cases. INSIGHT also demonstrates high prediction accuracies for simulation-intensive transient specifications given economic metric information. For example, using just DC Power information significantly improves INSIGHT’s prediction for Avg. Delay in comparators. Similarly, using just DC Power and Ratio information significantly improves INSIGHT’s prediction accuracy for Avg. Delay and Delay Balance. Table 5 presents INSIGHT’s prediction accuracy scores across 45nm, 130nm, and 180nm technologies when trained from scratch, using different dataset sizes. By employing simple transfer learning, we were able to reduce train dataset requirements successfully.
@@ 4.3. INSIGHT-M Performance
We evaluate the performance of INSIGHT-M against AutoCkt 17, an established online learning-based sizing algorithm based on PPO, using default hyperparameter settings. Table 6 presents the performance comparison for three different circuit types, OPAMP, TIA, and Comparator, on 45nm technology for three different test cases, each test case corresponding to a fixed target specification, randomly chosen from a set of designer-specified performance specifications. Performance is measured based on the number of real simulations (steps) required, with fewer steps indicating better performance (sample efficiency). The table also depicts the initial dataset size used to pre-train INSIGHT for each case. Without pre-training, the total number of steps (real-time simulator interactions) required falls within this mentioned pre-training dataset size for each case. As the table demonstrates, pre-trained INSIGHT enables INSIGHT-M to achieve convergence at only ¡ 20 real-time simulations across circuits for the given RL algorithm, achieving over 100-1000X improvement in sample efficiency over existing sizing methods.
@@ 5. Conclusion
This paper introduces INSIGHT, a novel, GPU-powered, technology-agnostic, effective universal neural simulator as a low-cost and high-fidelity substitute for standard circuit simulators in the analog design automation loop. INSIGHT facilitates effective and efficient design space exploration for improved sample efficiency. Leveraging self-attention to understand underlying inherent parameter-performance interdependencies within a given topology, INSIGHT accurately predicts performance at microsecond inference times across diverse circuit types and technologies. INSIGHT also enables accurate prediction of simulation-costly critical transient specifications using less expensive performance metric information. Our Model-Based Batch RL framework, INSIGHT-M, which utilizes INSIGHT for analog sizing, only requires ¡ 20 real-time simulations with 100-1000x lower simulation costs across circuits, showcasing the effectiveness of INSIGHT. To the best of our knowledge, this marks the first successful application of a universal neural simulator framework leveraging autoregressive Transformers for analog design automation. We also plan to extend this technology to address more complex analog design challenges, including mismatch analysis and technology design migration.